# Feature Request Analysis and Processing: Tasks, Techniques, and Trends

### v1

@@ 5.3. Approval Prediction
Early research employs machine learning classification models for automatic FRAP: Nizamani et al. (Nizamani et al., 2018) developed a Multinomial Naive Bayes (MNB) classifier, achieving an accuracy of 89.25%, outperforming models like Naive Bayes (NB), Support Vector Machine (SVM), Random Forest (RF), and LR. Umer et al. (Umer et al., 2019) enhanced this model by integrating sentiment analysis results from Senti4SD, which significantly improved predictive accuracy. Arshad et al. (Arshad et al., 2021) introduced a deep learning approach combining word2vec embeddings and sentiment analysis to capture syntactic and semantic information in feature requests, achieving superior performance compared to earlier methods by Nizamani et al. (Nizamani et al., 2018) and Umer et al. (Umer et al., 2019). Nafees et al. (Nafees and Rehman, 2021) compared SVM, MNB, and LR classifiers using a dataset of 40,000 enhancement reports from Bugzilla. Their study demonstrated that the SVM classifier outperformed MNB and LR, achieving an accuracy of 90.38% and an F-measure of 74.68%. This highlighted SVM’s effectiveness in handling high-dimensional feature spaces and identifying critical textual features.
Fitzgerald et al. (Fitzgerald et al., 2011, 2012) proposed predicting software faults by analyzing 13 distinct feature request attributes using Decision Trees (DT), NB, LR, and M5P-tree algorithms. Their findings highlighted that Bag-of-Words (BOW) and TF-IDF vectorization effectively predict fault-prone requests, emphasizing the value of textual content.
@@
Heppler et al. (Heppler et al., 2016) revealed that enhancement requests submitted by developers were twice as likely to succeed compared to external submissions. Building on this insight, Niu et al. (Niu et al., 2023) introduced a FRAP method that incorporates creator profiles (including ID, role as developer or not, and submission frequency), combined with TextCNN, which significantly enhances predictive performance. More recently, Zuo et al. (Zuo et al., 2025) improved FRAP accuracy by integrating generative LLM (e.g., GPT-4o, DeepSeek-V3) with creator profiles.
@@ 5.4.1. Lexical Analysis-Based Approaches
Figure 10. Common Software Feature Extraction Process
Table 5. Summary of Software Feature Extraction Research
Literature Feature Source Filtering Rules Application Year
[…]
(Bakiu and Guzman, 2017) User reviews Nouns, verbs, adjectives Software satisfaction analysis 2017
(Johann et al., 2017) App descriptions and user reviews POS patterns and sentence structures Software maintenance 2017
(Licorish et al., 2017) Application reviews Nouns Priority ranking 2017
[…]
Lexical analysis-based methods extract software features by identifying and filtering words or phrases that match predefined part-of-speech (POS) patterns—such as nouns, verbs, adjectives, or their combinations—to capture meaningful feature expressions from textual data. Based on these filtering rules, candidate software features are selected accordingly. Figure 10 illustrates the typical feature extraction workflow, while Table 5 summarizes the main sources and filtering rules employed in existing studies.
As shown in Table 5, the main sources of features are application reviews—including both expert reviews (Bakar et al., 2016) and user reviews (Bakiu and Guzman, 2017; Johann et al., 2017; Licorish et al., 2017; Dalpiaz and Parente, 2019)—and application descriptions (Johann et al., 2017). Filtering rules primarily rely on POS, often targeting specific POS combinations to identify meaningful feature phrases. The extracted features have been applied to various tasks such as requirement reuse (Bakar et al., 2016), user satisfaction analysis (Bakiu and Guzman, 2017), software maintenance (Johann et al., 2017), priority ranking (Licorish et al., 2017), and competitive product analysis (Dalpiaz and Parente, 2019).
Bakar et al. (Bakar et al., 2016) introduced an approach based on expert review texts, which preprocesses the data by selecting nouns, verbs, and adjectives, and then removes marginal documents through clustering. The method extracts feature phrases based on part-of-speech (POS) patterns, such as ¡adjective + noun¿ and ¡noun + adjective¿, capturing more representative feature phrases compared to manual extraction. Similarly, Bakiu et al. (Bakiu and Guzman, 2017) employed a collocation algorithm on user reviews, filtering for core POS tags and selecting features based on frequency and proximity, combined with lexical sentiment analysis to assess user satisfaction with the extracted features. Johann et al. (Johann et al., 2017) proposed the SAFE method, which relies on 18 manually defined POS patterns and five sentence structures to extract software features from both application descriptions and user reviews. Notably, SAFE requires minimal training data and has been widely adopted (Shah et al., 2019b, a).
Regarding priority ranking and competitive analysis, Licorish et al. (Licorish et al., 2017) focused on noun extraction from reviews to predict the processing priority of feature requests, while Dalpiaz et al. (Dalpiaz and Parente, 2019) utilized POS tagging and various POS combinations—such as ¡noun + adjective¿ and ¡adjective + verb¿—to mine features from competitor app reviews. This enabled the identification of strengths and weaknesses relative to competing products, thereby enhancing software competitiveness. Shah et al. (Shah et al., 2018) proposed annotation guidelines emphasizing the labeling of only consecutive word phrases containing nouns, limited to a maximum of three words, while excluding references to the application itself.
@@ 5.5. Intent Recognition
Early approaches combined NLP with fuzzy rule-based systems that analyze linguistic features at vocabulary, syntax, and semantic levels to identify sentence patterns corresponding to different intents such as explanation, advantage, disadvantage, and example. Integrating these fuzzy rules with classifiers such as random forests, naive Bayes, and support vector machines significantly improves classification performance (Shi et al., 2017). Building on this, tools like NERO (Mu et al., 2020) automate the classification of sentences within feature requests using such rule-based techniques. Beyond rule-based systems, methods grounded in speech-act theory classify user feedback at the sentence level into fine-grained intents, including requestive, suggestive, and assertive types. By aggregating these intent signals, developers gain clearer insights into user demands, aiding prioritization efforts (Kifetew et al., 2021). Decision tree–based classifiers have also been applied to detect specific intents, such as update requests in developer forums like Stack Overflow. Such studies reveal that a significant portion of user comments contain update requests, many of which receive timely responses, underscoring the practical importance of intent recognition in community feedback (Sheikhaei et al., 2023).
@@ 5.7. Classification
Table 6. Classification Categories in Existing Research
@@
(Li et al., 2018) Security, reliability, performance, life cycle, usability, ease of use, system interface KNN, NB, and SVM 2018
@@
(Tao et al., 2020) System, funds, spam, privacy, others Word co-occurrence matrix based on keywords 2020
@@
(Bai et al., 2024) Steal data, audio and video quality, customer service, hacker issues, meeting and account passwords, mute and unmute, features, office platform structural topic modeling with explainable AI 2023
@@ 5.9.1. Horizontal Traceability of Feature Requests
Due to the considerable effort involved in manually creating these horizontal relationships, existing methods predominantly rely on automated techniques leveraging text similarity. For instance, Natt et al. (Natt och Dag et al., 2004) implemented a cosine similarity-based vector space model to effectively link customer expectations with corresponding product requirements, significantly enhancing efficiency. Heck and Zaidman (Heck and Zaidman, 2014) employed TF-IDF and Latent Semantic Analysis (LSA) methods for text vectorization, demonstrating that TF-IDF notably outperformed LSA in identifying novel feature request connections. Furthermore, Merten et al. (Merten et al., 2016b) evaluated various information retrieval algorithms, including TF-IDF, LSI, and variants of BM25, concluding that prioritizing feature request titles over comments substantially improves linking accuracy.
@@ 5.10. Developer Assignment
Once a feature request is approved, a critical step is assigning it to a suitable developer. Due to variation in developer expertise, proper assignment plays a vital role in reducing implementation failures and improving development efficiency (Mao et al., 2015). Most developer assignment techniques are based on the assumption that ··either the original author of the code or a developer with similar prior experience is the best candidate for the task.”
Developer assignment approaches can generally be categorized into three groups:
@@
✎ Artifact-Centric Approaches. These methods first retrieve source code artifacts (e.g., files, methods, classes) related to the feature request, and then identify developers associated with those artifacts. Kagdi et al. (Kagdi and Poshyvanyk, 2009; Kagdi et al., 2012) created a document index of code units using identifiers and comments, and employed Latent Semantic Indexing (LSI) to match requests with code components. Developers were ranked based on their contribution distance, calculated using the Euclidean distance between commit vectors. Linares-Vásquez et al. (Linares-Vásquez et al., 2012) improved accessibility by avoiding repository mining, instead linking requests to authors extracted from annotated source files using srcML (Collard et al., 2013). Hossen et al. (Hossen et al., 2014) introduced the iMacPro method, which ranks relevant code components by historical modification frequency and extracts both maintainers and original authors for recommendations. It demonstrated superior recall compared to earlier methods, such as iA (Linares-Vásquez et al., 2012) and iAcPro (Hossen et al., 2014). Zanjani et al. (Zanjani et al., 2015) proposed iHDev, which used TF-IDF and KNN to identify relevant code units, then used contribution history to rank developers. It outperformed xFinder (Kagdi et al., 2012) and iMacPro (Hossen et al., 2014) in both recall and Mean Reciprocal Rank (MRR).
@@ 5.11. Code Localization
Palomba et al. (Palomba et al., 2017) proposed CHANGEADVISOR, a method that leverages user feedback to support code localization. It begins by analyzing user comments using the ARdoc (Panichella et al., 2016) technique, which captures sentence structure, semantics, and sentiment. The feedback is clustered using topic modeling techniques such as LDA, LDA-GA (which combines LDA with genetic algorithms), and HDP (Hierarchical Dirichlet Process). To associate feedback clusters with code classes, they compute asymmetric Dice similarity scores. Code elements that exceed a similarity threshold are ranked as likely targets for implementation. The method demonstrated high accuracy in identifying code impacted by user-requested changes and outperformed baseline techniques in both precision and recall. Ciurumelea et al. (Ciurumelea et al., 2017) introduced the URR approach, which analyzes user reviews to identify code segments needing modification. They apply Gradient Boosted Regression Trees (GBRT) to classify reviews into six high-level and twelve fine-grained categories. Code artifacts (files, classes, and methods) are then annotated with structural types and indexed accordingly. Using Apache Lucene, the system performs a relevance search between the processed review content and the indexed source files. The most relevant files are recommended as targets for change, demonstrating an effective mapping between high-level user concerns and low-level implementation details. Kato and Goto (Kato and Goto, 2017) proposed an innovative interaction model within a web-based Integrated Development Environment (IDE), called User-Generated Variables (UGV). This approach enables end users to specify which parameters they wish to adjust in the application—without needing to read or understand the underlying source code directly. These UGVs are automatically linked to corresponding variable declarations within the codebase, streamlining the code localization process for non-technical stakeholders. This method shifts the burden of interpretation from users to the tooling, enhancing accessibility and usability in feature-driven development contexts.
@@ 5.12. Code Recommendation
6. Publicly available dataset and tools
@@ 6.2. Publicly Available Datasets
A significant impediment in many areas of requirements engineering research is the scarcity of publicly accessible datasets, often due to the proprietary nature of software requirements. However, the domain of feature requests presents a unique advantage: a substantial volume of data originates from open platforms, such as public repositories and app stores. This accessibility, often leveraged through web scraping techniques, has partially alleviated the data scarcity challenge that plagues other RE sub-fields. Consequently, various datasets have been curated and made publicly available by researchers. Table LABEL:tab:dataset_long catalogues these resources, detailing their origins and the specific research tasks they are typically used for (e.g., training classification models, evaluating duplicate detection).
@@
(Zhang et al., 2019) https://github.com/heulaoyoutiao/bugtag2 2019

### v2

@@ 5.3. Approval Prediction
Early research employs machine learning classification models for automatic FRAP: Nizamani et al. (Nizamani et al., 2018) developed a MNB classifier, achieving an accuracy of 89.25%, outperforming models like NB, SVM, RF, and LR. Umer et al. (Umer et al., 2019) enhanced this model by integrating sentiment analysis results from Senti4SD, which significantly improved predictive accuracy. Arshad et al. (Arshad et al., 2021) introduced a deep learning approach combining word2vec embeddings and sentiment analysis to capture syntactic and semantic information in feature requests, achieving superior performance compared to earlier methods by Nizamani et al. (Nizamani et al., 2018) and Umer et al. (Umer et al., 2019). Nafees et al. (Nafees and Rehman, 2021) compared SVM, MNB, and LR classifiers using a dataset of 40,000 enhancement reports from Bugzilla. Their study demonstrated that the SVM classifier outperformed MNB and LR, achieving an accuracy of 90.38% and an F-measure of 74.68%. This highlighted SVM’s effectiveness in handling high-dimensional feature spaces and identifying critical textual features.
Fitzgerald et al. (Fitzgerald et al., 2011, 2012) proposed predicting software faults by analyzing 13 distinct feature request attributes using DT, NB, LR, and M5P-tree algorithms. Their findings highlighted that BOW and TF-IDF vectorization effectively predict fault-prone requests, emphasizing the value of textual content.
@@
Heppler et al. (Heppler et al., 2016) revealed that enhancement requests submitted by developers were twice as likely to succeed compared to external submissions. Building on this insight, Niu et al. (Niu et al., 2025a) introduced a FRAP method that incorporates creator profiles (including ID, role as developer or not, and submission frequency), combined with TextCNN, which significantly enhances predictive performance. More recently, Zuo et al. (Zuo et al., 2025) improved FRAP accuracy by integrating generative LLM (e.g., GPT-4o, DeepSeek-V3) with creator profiles.
@@ 5.4.1. Lexical Analysis-Based Approaches
Lexical analysis-based methods extract software features by identifying and filtering words or phrases that match predefined POS patterns, such as nouns, verbs, adjectives, or their combinations, to capture meaningful feature expressions from textual data. Based on these filtering rules, candidate software features are selected accordingly.
Bakar et al. (Bakar et al., 2016) introduced an approach based on expert review texts, which preprocesses the data by selecting nouns, verbs, and adjectives, and then removes marginal documents through clustering. Similarly, Bakiu et al. (Bakiu and Guzman, 2017) employed a collocation algorithm on user reviews, filtering for core POS tags and selecting features based on frequency and proximity, combined with lexical sentiment analysis to assess user satisfaction with the extracted features. Johann et al. (Johann et al., 2017) proposed the SAFE method, which relies on 18 manually defined POS patterns and five sentence structures to extract software features from both application descriptions and user reviews. Notably, SAFE requires minimal training data and has been widely adopted (Shah et al., 2019b, a).
Regarding priority ranking and competitive analysis, Licorish et al. (Licorish et al., 2017) focused on noun extraction from reviews to predict the processing priority of feature requests, while Dalpiaz et al. (Dalpiaz and Parente, 2019) utilized POS tagging and various POS combinations—such as ¡noun + adjective¿ and ¡adjective + verb¿—to mine features from competitor app reviews. Shah et al. (Shah et al., 2018) proposed annotation guidelines emphasizing the labeling of only consecutive word phrases containing nouns, limited to a maximum of three words, while excluding references to the application itself.
@@ 5.5. Intent Recognition
Early approaches combined NLP with fuzzy rule-based systems that analyze linguistic features at vocabulary, syntax, and semantic levels to identify sentence patterns corresponding to different intents such as explanation, advantage, disadvantage, and example. Integrating these fuzzy rules with classifiers such as RF, NB, and SVM significantly improves classification performance (Shi et al., 2017). Building on this, tools like NERO (Mu et al., 2020) automate the classification of sentences within feature requests using such rule-based techniques. Beyond rule-based systems, methods grounded in speech-act theory classify user feedback at the sentence level into fine-grained intents, including requestive, suggestive, and assertive types. By aggregating these intent signals, developers gain clearer insights into user demands, aiding prioritization efforts (Kifetew et al., 2021). DT–based classifiers have also been applied to detect specific intents, such as update requests in developer forums like Stack Overflow. Such studies reveal that a significant portion of user comments contain update requests, many of which receive timely responses, underscoring the practical importance of intent recognition in community feedback (Sheikhaei et al., 2023).
@@ 5.7. Classification
Table 4. Classification Categories in Existing Research.
@@
[…]
@@
[…]
@@
[…]
@@ 5.9.1. Horizontal Traceability of Feature Requests
Due to the considerable effort involved in manually creating these horizontal relationships, existing methods predominantly rely on automated techniques leveraging text similarity. For instance, Natt et al. (Natt och Dag et al., 2004) implemented a cosine similarity-based vector space model to effectively link customer expectations with corresponding product requirements, significantly enhancing efficiency. Heck and Zaidman (Heck and Zaidman, 2014) employed TF-IDF and Latent Semantic Analysis (LSA) methods for text vectorization, demonstrating that TF-IDF notably outperformed LSA in identifying novel feature request connections. Furthermore, Merten et al. (Merten et al., 2016b) evaluated various information retrieval algorithms, including TF-IDF, Latent Semantic Indexing (LSI), and variants of BM25, concluding that prioritizing feature request titles over comments substantially improves linking accuracy.
@@ 5.10. Developer Assignment
Once a feature request is approved, a critical step is assigning it to a suitable developer. Due to variation in developer expertise, proper assignment plays a vital role in reducing implementation failures and improving development efficiency (Mao et al., 2015). Most developer assignment techniques are based on the assumption that “either the original author of the code or a developer with similar prior experience is the best candidate for the task.”
Developer assignment is conceptually rooted in two long-standing research streams that address the same core challenge of matching a natural-language issue description to an appropriate expert: bug report assignment and ticket routing.
Bug report assignment originated from IR-based lexical matching between bug reports and developer artifacts (Matter et al., 2009; Anvik and Murphy, 2011). A seminal study by Anvik et al. (Anvik et al., 2006) first formulated bug report assignment as a supervised text classification problem, training models such as NB and SVM on historical project data. Each report was represented using textual and categorical features (e.g., summary, description, and component). The trained models predicted the most likely developers for new reports. Over the years, techniques have gradually evolved toward data-driven and semantic reasoning, incorporating historical fixes, severity, and activity traces into ranking-based and deep neural models (Xuan et al., 2014; Tian et al., 2016; Lee et al., 2017; Mani et al., 2019). Recent advances further leverage scheduling and dependency constraints, contrastive objectives, and large-scale industrial evaluations, transforming bug triage into a robust, deployable automation process (Jahanshahi and Cevik, 2022; Borg et al., 2024; Wang et al., 2024).
Ticket routing, emerging from enterprise service management, advanced from probabilistic and sequence-mining models for routing efficiency (Shao et al., 2008b, a) to graph and embedding-based expert representations that capture both technical competence and organizational relations (Zhou et al., 2016; Xu and He, 2018; Mandal et al., 2019). More recent systems leverage LLMs to jointly predict responsible teams and resolvers while exploiting organizational knowledge for retrieval-augmented and relation-aware routing (Feng et al., 2022; Zangari et al., 2023; Liu et al., 2025).
Together, these developments provide the conceptual and methodological groundwork for contemporary research on feature request assignment, where the focus shifts from fixing to implementing new functionality while preserving the same principle of expertise-oriented task routing.
Feature request assignment approaches can generally be categorized into three groups:
@@
✎ Artifact-Centric Approaches. These methods first retrieve source code artifacts (e.g., files, methods, classes) related to the feature request, and then identify developers associated with those artifacts. Kagdi et al. (Kagdi and Poshyvanyk, 2009; Kagdi et al., 2012) created a document index of code units using identifiers and comments, and employed LSI to match requests with code components. Developers were ranked based on their contribution distance, calculated using the Euclidean distance between commit vectors. Linares-Vásquez et al. (Linares-Vásquez et al., 2012) improved accessibility by avoiding repository mining, instead linking requests to authors extracted from annotated source files using srcML (Collard et al., 2013). Hossen et al. (Hossen et al., 2014) introduced the iMacPro method, which ranks relevant code components by historical modification frequency and extracts both maintainers and original authors for recommendations. It demonstrated superior recall compared to earlier methods, such as iA (Linares-Vásquez et al., 2012) and iAcPro (Hossen et al., 2014). Zanjani et al. (Zanjani et al., 2015) proposed iHDev, which used TF-IDF and KNN to identify relevant code units, then used contribution history to rank developers. It outperformed xFinder (Kagdi et al., 2012) and iMacPro (Hossen et al., 2014) in both recall and Mean Reciprocal Rank (MRR).
@@ 5.11. Code Localization
One of the earliest and most influential studies in this field is by Marcus et al. (Marcus et al., 2004), who introduced an IR-based approach using LSI for concept and feature localization. The method represents source code entities as textual documents and computes their semantic similarity with feature request descriptions in a latent space, enabling the identification of related code even without exact keyword matches. This seminal work laid the foundation for later IR-based and hybrid localization techniques.
Recent surveys have shown that IR-based methods in software maintenance have progressively evolved from surface-level textual retrieval to semantically enriched and structure-aware reasoning (Niu et al., 2025b). Feature location, a closely related and historically rich research area, has evolved through several major phases over the past two decades. Early studies conceptualized feature location as an IR problem, where natural-language descriptions of software features were matched to relevant source code entities based on textual similarity (Xue et al., 2012; Hoseini et al., 2014; Binkley et al., 2015). Subsequent research moved beyond pure textual analysis to incorporate structural, historical, and dynamic information, resulting in hybrid approaches that combined static dependencies, execution traces, and change histories to improve localization accuracy (Peng et al., 2013; Chochlov et al., 2017; Font et al., 2017). With the rise of deep learning, feature location began to adopt semantic representation learning techniques, replacing manually engineered IR features with distributed embeddings that better captured conceptual relationships between feature descriptions and source code (Corley et al., 2015; Razzaq et al., 2020; Pérez et al., 2020). More recently, the field has advanced toward intelligent and hybrid feature location, integrating semantic, structural, and behavioral signals as well as traceability and feature characteristics to achieve more robust and context-aware localization (Echeverría et al., 2021; Razzaq et al., 2021; Mukelabai et al., 2023). This long-standing evolution, from IR-based retrieval to semantically enriched and hybrid reasoning, has directly shaped modern code localization research.
Building on this foundation, Palomba et al. (Palomba et al., 2017) proposed CHANGEADVISOR, which leverages user feedback for code localization. It analyzes comments with ARdoc (Panichella et al., 2016) to capture structure, semantics, and sentiment, then clusters feedback via LDA, LDA-GA, and HDP. Feedback clusters are linked to code classes using asymmetric Dice similarity; elements exceeding a threshold are ranked as likely implementation targets. The method achieves high accuracy and outperforms baselines in both precision and recall. Where2Change (Zhang et al., 2019a) advanced CHANGEADVISOR by introducing Word2Vec to enrich semantics and applying a weight-selection cosine similarity to rank code classes, yielding higher Top-N accuracy and recall across 10 open-source apps. Furthermore, RISING (Zhou et al., 2020) leverages commit messages to bridge the vocabulary gap between reviews and source files, leading to better localization. Hu et al. (Hu et al., 2022) introduced a benchmark for user-review–based code localization: they curated gold-standard $\langle\text{review},\text{source file}\rangle$ links for mobile apps and compared supervised with unsupervised approaches, with GraphCodeBERT consistently outperforming the baselines. Ciurumelea et al. (Ciurumelea et al., 2017) propose URR, which analyzes user reviews to localize change. Reviews are classified via Gradient Boosted Regression Trees into six high-level and twelve fine-grained categories. Code artifacts (files/classes/methods) are annotated and indexed, and Apache Lucene retrieves relevant files by matching processed reviews to the index, effectively mapping user concerns to implementation targets. Kato and Goto (Kato and Goto, 2017) propose User-Generated Variables (UGV), a web-based IDE interaction model. UGV lets end users specify adjustable parameters without reading code; the system auto-links them to variable declarations, streamlining localization for non-technical stakeholders. This shifts interpretation to tooling, improving accessibility in feature-driven development. Recently, YourCoLo (Chi et al., [n. d.]) is a user review-based code localization approach that models one-to-many review-to-code relationships and incorporates inter-code structural connections. It leverages project-level contextual signals and a customized ranking mechanism to more accurately identify relevant code elements linked to user feedback, outperforming SOTA baselines.
Similar to code localization for feature requests, IR-based bug localization (IRBL) (Zhou et al., 2012; Niu et al., 2025b) aims to identify source code implicated by bug reports. Historical feature requests have been shown to provide valuable contextual signals for IRBL approaches (Niu et al., 2023, 2024; Rath et al., 2018; Yu et al., 2022). Likewise, Xiao et al. (Xiao et al., 2023) leveraged bug report information to enhance code localization for user reviews. Maarleveld et al. (Maarleveld et al., 2025) empirically examined the performance of IR-based file localization for issue reports (including both bug reports and feature requests), emphasizing the importance of models tailored to general issue localization and project-specific characteristics.
Recently, NoCode-bench (Deng et al., 2025) benchmarks LLMs on natural-language-driven feature addition, bridging feature requests and code localization. The dataset contains 634 real tasks from 10 projects, pairing documentation updates with multi-file code changes. Nine state-of-the-art generative LLMs were evaluated, revealing their limited ability to map natural-language feature descriptions to the correct code regions. Besides, SWE-Bench (Jimenez et al., 2023) and its extensions (including SWE-Bench Lite (Jimenez et al., 2023), SWE-Bench Multimodal (Yang et al., 2024b), and SWE-Bench Verified (Chowdhury et al., 2024)) benchmark real-world issue resolution, encompassing both bug reports and feature requests. A series of LLM-based approaches (Reddy et al., 2025; Xia et al., 2024; Chen et al., 2025; Yang et al., 2024a) have been evaluated on these benchmarks to assess their capability in automating issue resolution tasks.
@@ 5.12. Code Recommendation
6. Publicly Available Datasets and Tools
@@ 6.2. Publicly Available Datasets
A significant impediment in many areas of requirements engineering research is the scarcity of publicly accessible datasets, often due to the proprietary nature of software requirements (Yu et al., 2024). However, the domain of feature requests presents a unique advantage: a substantial volume of data originates from open platforms, such as public repositories and app stores. This accessibility, often leveraged through web scraping techniques, has partially alleviated the data scarcity challenge that plagues other RE sub-fields. Consequently, various datasets have been curated and made publicly available by researchers. Table LABEL:tab:dataset_long catalogs these resources, detailing their origins and the specific research tasks they are typically used for (e.g., training classification models, evaluating duplicate detection).
@@
(Zhang et al., 2019b) https://github.com/heulaoyoutiao/bugtag2 2019
@@
(Hu et al., 2022) https://github.com/lcynju/review2code Feature localization. 2022